I did check the other questions and some were similiar but not quite.

It sounds easy but I just don't know where to start.

Answer by malakumar_kos@yahoo.com(315)

Two cars leave an intersection at the same time. One travels north and the other heads west traveling 10mph faster. After 1 hour, they are 50 miles apart. What were their speeds?
Let the speed of the car travelling towards north be = x m/hr
then the speed of the car travelling towards west = (x+10)m/hr
after 1 hr distance between them is given to be = 50 miles
After 1 hr distance travelled by the car towards north = x.1 = x miles
After 1 hr distance travelled by the car towards west = (x+10).1 = (x+10) mil
therefore using pythagoras theorem we get (50)^2 = x^2+(x+10)^2
2500 = x^2+x^2+20x+100
2500-100 = 2x^2+20x
2400 = 2x^2+20x
dividing by 2 throughout
we get 1200 = x^2+10x
x^2+10x-1200 = 0
x^2+40x-30x-1200=0
x(x+40)-30(x+40) = 0
(x+40).(x-30) = 0 or x-30 = 0 or x = 30
speed of the car towards north is = 30 miles/hr
speed of the car towards west is = 40 miles/hr